OpenStudy (anonymous):

Solve the system using the addition method. 5x - 2y = 3 2x - y = 0 What is the y-value in the solution? 3 6 -3 none of the above

OpenStudy (anonymous):

\[5x-2y=3\] equation 1 \[2x-y=0\] equation 2 by applying on equation 2 \[\therefore y=2x\] by substituting in equation 1 \[\therefore 5x-2(2x)=3\] \[\therefore 5x-4x=3\] \[\therefore x=3\] so by substituting in any equation u get (y=6)

OpenStudy (anonymous):

love u

OpenStudy (anonymous):

thnx :$ any other question ?